Two objects are located along the y-axis. Object A has a charge of +30.0 µC and is located at the origin. Object B has a charge of -22.0 µC and is located a distance of 8.00 cm above object A. Determine the force between these two objects. 



1
Expert's answer
2021-01-01T14:07:08-0500

"F=k\frac{|q_1||q_2|}{r^2}=9\cdot10^{9}\cdot\frac{30\cdot10^{-6}\cdot22\cdot10^{-6}}{0.08^2}=928(N)" .